LOG-IN
Displaying reports 1-20 of 12923.Go to page 1 2 3 4 5 6 7 8 9 10 End
MIF (General)
Print this report.
masayuki.nakano - 19:51, Monday 25 May 2020 (14421)Get code to link to this report
IFO recovery 0525

The IFO has been recovered. 

  • LSC guardian could reach to OBSERVATION
  • Due to the power drop of the GRY, the FNC cannot been locked. Therefore, I and YamaT-sensei enter the mine and align the green fiber as suggested by Sugimoto-kun 
  • I changed the FNC gain to -20 dB and GRY-PDH gain to 6. The UGF of PDH loop is around 4 kHz.
  • YamaT-sensei fixed the IX Tcam.
  • The alignment is terrible. I guess we need to tweak the IMMTs.
ChiefMeeting (General)
Print this report.
shinji.miyoki - 17:11, Monday 25 May 2020 (14420)Get code to link to this report
Chief Meeting 200525

[REPORT]

0.Chief Meeting

  • Subgroup reports are necessary? 
  • check interface with each subgroup.
  • limit to chief, sub-chief?
  • SEO will decide waht is chief meeting for upgrade.
  • New wiki system will come. --> We move to this new page. Target is in this FY.

1. Schedule of After Observation

  • Suspension Mode from April 27th to May 31st
    • KAGRA was in  "Suspended Mode".
    • UT keeps level 3 until May 31st. We obey this rule. The next possible level change might be around 1st June.
    • In this mode, we keep IFO just in "DOWN state" in the LSC_GUARDIAN to be able to quickly recover the IFO to PRFPMI. Most important thing is to keep arm GR resonances in both arms by daily shift members.
    • Because of so many earthquakes, especially around Kamikochi area, we decided to concentrate in keeping IFO healthy and safe. However, some remote/in control room tasks are allowed at present because of better COVID-19 situation. However, we should prepare for one or more  M5 class earthquakes in the future according to earthquake specialists in Nagoya-U.
    • Some members are now making a program to change Type-A state automatically to damping mode according to the amount of shaking due to earthquakes.
    • The shift schedule is in here
  • Post Observation Tasks
    • Assuming the start of 1st June, more activities are expected. Let's make schedule.
    • from 9 to 17 working time is default. (Only tasks that much less possibilities to call several experts might be approved from 17~22 in the control room/remote. <- On discussion in SEO)
    • Keep remote the task rule as defined before.
  • O3c or O4 schedule
    • Decision for O3c in Virgo will be done within two weeks (from EO). We don't know what is going in LIGO.
    • We keep "post obs tasks" because it dose not depend on O3c / O4.
    • We will prepare for upgrade tasks assuming no O3c before Virgo/LIGO decision about O3c and its length. If decision will be done, we reconsider KAGRA side schedule and what we should do.
    • In O3c case (to be discussed after)
      • We confirmed that KAGRA will join O3c if LV go O3c and this message has been decleared by PI to LV. One of reason is because it will be "good" (experience etc) for "KAGRA collaborator".
      • O3c restart will seem to require 4 or more weeks preparation including ER. 
      • The actual obaservation length for O3c by KAGRA will (might) be discussed in EO accounting for actual KAGRA scientific contribution, recovery of relative delay from LV, etc. 
      • What KAGAR can prepar for O3c --> Discussion.
    • In O4 case
      • Decide major upgrade items and their schedule (almost done with delay due to COVID-19) --> Discussion.
      • Select possible/desirable tasks among additionally proposed task lits. --> Discussion.

2. JRPC related

  • No JRPC meeting for the last rwo weeks.
  • CAL homeworks status
    •  
    • The next review will be on May 31st. 

 

3. KSC Matters

  • LV white paper will be made. KAGRA side managers were selected to Tagoshi-san, and Miyakawa-kun.
    • LSC will announce, after that KAGRA will ask you ( limit is June 22nd)
  • Schedule
    • The KAGRA telecon, May 27 (Wed) & 28 (Thur), 2020.
    • The 25th KAGRA Face-to-Face meetingat U. Toyama, August 20-22, 2020.
    • The LVK collaboration meeting at Cardiff, UK, September 14-17, 2020 (JPS meeting has same schedule) ---> they hold it in remote style is under discussion.
    • The 26th KAGRA Face-to-Face meetingat NCU, Taoyuan City, Taiwan, Dec. 17-18, 2020.
    • The 7th KAGRA International Workshop in NCU, Taoyuan City, Taiwan, Dec. 18-20, 2020
    • The LVK collaboration meeting,Lake Geneva, Wisconsin, USA, March 22-26 ??, 2021
    • The 8th KAGRA International Workshop, in Daejeon, Korea, sometime, 2021
    • The 9th KAGRA International Workshop, in Beijing, China, sometime, 2021
    • Other related GW meetings are listed in http://gwwiki.icrr.u-tokyo.ac.jp/JGWwiki/KAGRA/DAWG/conferences

4. Papers

  • A plan is summarized in here . The author and/or manager were almost fixed.
  • PTEP2 noise budget paper is managed by K.Yamamoto-kun (Toyama-U)

 

5. Others

  • Application for presentation for JPS (on-line) has started. Please each chief inform your group possible presentations to Uchiyama-kun.
  • Observation Talks are necesary.

 

[Discussions]

1. Post Observation task plan and scheduling (Miyakawa) in June

  • Proposed plans. 
  • We should prioratize them and make a schedule.
  • At least two weeks can be used for this.

2.  Waht we can prepare for O3c (in level2)

  • We expect only small amount of improvements such as,
    • OMC stray light check, -> 1 week (5 days) ? 
    • PD for OMC change, -> 1 week? (3 weeks ?) --> OMC opening tasks.
    • DRFPMI / ASC trial if enough weeks can be adopted. 2 weeks (10 days) ?
    • FFU operation and temp control. Particle counter check -> PEM 
    • SRM/2/3 suspension update is necessary if DRFPMI (1 week )
      • necessary for OMC task. 
    • Sensor correction for SRs by IXV seismometers (software change). 
  • Drastic sensitivity enhancement from around 1 Mpc cannot be expected because almost noise reduction especially below 200 Hz requires actions for O4.

 

5. Schedule for O4 and any catch up plan ?

  • Plan (but no delay case)
  • Additional Plan (budget allocation is not guaranteed)
    • Offline known line noise subtraction. (PEM / DAS / DGS / DMG / CAL)
    • Scattered light around IOO (AOS: Akutsu)
    • IMMT2 suspension frame modification to keep a clear beam path of transmitting IR and oplev (AOS, VIS). No document about layout.
    • BS, PRs, and ETMXY surface monitors (VAC, AOS)
    • Electrical ground optimization (FCL / PEM).
    • TMS P/Y coupling (AOS or CRY). Firstly, we should identify its reason.
    • High power laser pre-preparation (IOO + Haino-san). We might not use it, however we will put in PSL room for swapping in the future. (IOO)
    • DGS update will be done just after post-observation tasks. (DGS)
    • Additional GVs in MICH area for more convenient vacuum opening task. (VAC)
    • More QPDs for monitoring on POP tables. (IOO/MIF)
    • Interlock preparation. (Aso)
    • Poor lifetime cryocoolers replacement. (CRY)
    • In vacuum PD preparation for ASC/REFL. (MIF)
    • In vacuum intensity noise reduction system plan. The technical design meeting will be managed by Moriwaki-sensei. (IOO)
    • Low loss mirror preparation for OMC. (IOO)
    • DGS/AEL upgrade. (DGS/AEL)
    • Microseismic noise compensator. (TBD) -> Make team. 
    • VAC monitor

 

[ Past Chief Meeting Page ] will be closed because of very poor security and no more maintenacne. Other machine will be discussed as servers.

ChiefMeeting (General)
Print this report.
yuki.inoue - 15:22, Monday 25 May 2020 (14419)Get code to link to this report
Comment to May 11th (Click here to view original report: 14364)
CAL report:
https://gwdoc.icrr.u-tokyo.ac.jp/cgi-bin/private/DocDB/ShowDocument?docid=11674
ChiefMeeting (General)
Print this report.
shinji.miyoki - 18:38, Saturday 23 May 2020 (14418)Get code to link to this report
Chief Meeting 200525

[REPORT]

0.Chief Meeting

  • Subgroup reports are necessary? 
  • check interface with each subgroup.
  • limit to chief, sub-chief?
  • SEO will decide waht is chief meeting for upgrade.
  • New wiki system will come. --> We move to this new page. Target is in this FY.

1. Schedule of After Observation

  • Suspension Mode from April 27th to May 31st
    • KAGRA was in  "Suspended Mode".
    • UT keeps level 3 until May 31st. We obey this rule. The next possible level change might be around 1st June.
    • In this mode, we keep IFO just in "DOWN state" in the LSC_GUARDIAN to be able to quickly recover the IFO to PRFPMI. Most important thing is to keep arm GR resonances in both arms by daily shift members.
    • Because of so many earthquakes, especially around Kamikochi area, we decided to concentrate in keeping IFO healthy and safe. However, some remote/in control room tasks are allowed at present because of better COVID-19 situation. However, we should prepare for one or more  M5 class earthquakes in the future according to earthquake specialists in Nagoya-U.
    • Some members are now making a program to change Type-A state automatically to damping mode according to the amount of shaking due to earthquakes.
    • The shift schedule is in here
  • Post Observation Tasks
    • Assuming the start of 1st June, more activities are expected. Let's make schedule.
    • from 9 to 17 working time is default. (Only tasks that much less possibilities to call several experts might be approved from 17~22 in the control room/remote. <- On discussion in SEO)
    • Keep remote the task rule as defined before.
  • O3c or O4 schedule
    • Decision for O3c in Virgo will be done within two weeks (from EO). We don't know what is going in LIGO.
    • We keep "post obs tasks" because it dose not depend on O3c / O4.
    • We will prepare for upgrade tasks assuming no O3c before Virgo/LIGO decision about O3c and its length. If decision will be done, we reconsider KAGRA side schedule and what we should do.
    • In O3c case (to be discussed after)
      • We confirmed that KAGRA will join O3c if LV go O3c and this message has been decleared by PI to LV. One of reason is because it will be "good" (experience etc) for "KAGRA collaborator".
      • O3c restart will seem to require 4 or more weeks preparation including ER. 
      • The actual obaservation length for O3c by KAGRA will (might) be discussed in EO accounting for actual KAGRA scientific contribution, recovery of relative delay from LV, etc. 
      • What KAGAR can prepar for O3c --> Discussion.
    • In O4 case
      • Decide major upgrade items and their schedule (almost done with delay due to COVID-19) --> Discussion.
      • Select possible/desirable tasks among additionally proposed task lits. --> Discussion.

2. JRPC related

  • No JRPC meeting for the last rwo weeks.
  • CAL homeworks status
    •  
    • The next review will be on May 31st. 

 

3. KSC Matters

  • LV white paper will be made. KAGRA side managers were selected to Tagoshi-san, and Miyakawa-kun.
  • Schedule
    • The KAGRA telecon, May 27 (Wed) & 28 (Thur), 2020.
    • The 25th KAGRA Face-to-Face meetingat U. Toyama, August 20-22, 2020.
    • The LVK collaboration meeting at Cardiff, UK, September 14-17, 2020 (JPS meeting has same schedule) ---> they hold it in remote style is under discussion.
    • The 26th KAGRA Face-to-Face meetingat NCU, Taoyuan City, Taiwan, Dec. 17-18, 2020.
    • The 7th KAGRA International Workshop in NCU, Taoyuan City, Taiwan, Dec. 18-20, 2020
    • The LVK collaboration meeting,Lake Geneva, Wisconsin, USA, March 22-26 ??, 2021
    • The 8th KAGRA International Workshop, in Daejeon, Korea, sometime, 2021
    • The 9th KAGRA International Workshop, in Beijing, China, sometime, 2021
    • Other related GW meetings are listed in http://gwwiki.icrr.u-tokyo.ac.jp/JGWwiki/KAGRA/DAWG/conferences

4. Papers

  • A plan is summarized in here . The author and/or manager were almost fixed.
  • PTEP2 noise budget paper is managed by K.Yamamoto-kun (Toyama-U)

 

5. Others

  • Application for presentation for JPS (on-line) has started. Please each chief inform your group possible presentations to Uchiyama-kun.

 

 

[Discussions]

1. Post Observation task plan and scheduling (Miyakawa) in June

  • Proposed plans. 
  • We should prioratize them and make a schedule.
  • At least two weeks can be used for this.

2.  Waht we can prepare for O3c

  • We expect only small amount of improvements such as,
    • OMC stray light check, -> 1 week (5 days) ?
    • PD for OMC change, -> 1 week?
    • DRFPMI / ASC trial if enough weeks can be adopted. 2 week (10 days) ?
    • ??
  • Drastic sensitivity enhancement from around 1 Mpc cannot be expected because almost noise reduction especially below 200 Hz requires actions for O4.

 

5. Schedule for O4 and any catch up plan ?

  • Plan (but no delay case)
  • Additional Plan (budget allocation is not guaranteed)
    • Offline known line noise subtraction. (PEM / DAS / DGS / DMG / CAL)
    • Scattered light around IOO (AOS)
    • IMMT2 suspension frame modification to keep a clear beam path of transmitting IR and oplev (AOS, VIS). No document about layout.
    • BS, PRs, and ETMXY surface monitors (VAC, AOS)
    • Electrical ground optimization (FCL / PEM).
    • TMS P/Y coupling (AOS or CRY). Firstly, we should identify its reason.
    • High power laser pre-preparation (IOO + Haino-san). We might not use it, however we will put in PSL room for swapping in the future. (IOO)
    • DGS update will be done just after post-observation tasks. (DGS)
    • Additional GVs in MICH area for more convenient vacuum opening task. (VAC)
    • More QPDs for monitoring on POP tables. (IOO/MIF)
    • Interlock preparation. (Aso)
    • Poor lifetime cryocoolers replacement. (CRY)
    • In vacuum PD preparation for ASC/REFL. (MIF)
    • In vacuum intensity noise reduction system plan. The technical design meeting will be managed by Moriwaki-sensei. (IOO)
    • Low loss mirror preparation for OMC. (IOO)
    • DGS/AEL upgrade. (DGS/AEL)
    • Microseismic noise compensator. (TBD)

 

[ Past Chief Meeting Page ] will be closed because of very poor security and no more maintenacne. Other machine will be discussed as servers.

MIF (General)
Print this report.
kouseki.miyo - 21:09, Friday 22 May 2020 (14416)Get code to link to this report
Comment to On site shift report (Click here to view original report: 14414)

report as a remote shift
Trans powers of both arm are OK. (~0.7/X, ~0.7/Y)
IMC is OK. (2.3W)

VIS (SR2)
Print this report.
fabian.arellano - 15:37, Friday 22 May 2020 (14415)Get code to link to this report
Comment to Decay time measurements. (Click here to view original report: 14413)

Today I continued with the measurements.

Pending review of the quality of the data which may suggest repeating a measurement, I have finished the measurements with the control system on.

I still need to do the measurements with the control system off.

MIF (General)
Print this report.
fabian.arellano - 15:34, Friday 22 May 2020 (14414)Get code to link to this report
On site shift report

The amount of green light transmitted in the arms is

  • Y Arm: 0.7 (0.5 minimum allowed).
  • X Arm: 0.8 (0.5 minimum allowed).

The IMC output power is 2.3 W.

According to the guidelines, these values mean that the orientation of the mirrors don't deserve adjustment.

Comments related to this report
kouseki.miyo - 21:09, Friday 22 May 2020 (14416)

report as a remote shift
Trans powers of both arm are OK. (~0.7/X, ~0.7/Y)
IMC is OK. (2.3W)

VIS (SR2)
Print this report.
fabian.arellano - 18:14, Thursday 21 May 2020 (14413)Get code to link to this report
Decay time measurements.

I began measuring the decay times of the supension.

The files are in the following directory:

/kagra/Dropbox/Subsystems/VIS/TypeBData/SR2/quality_factors/

Comments related to this report
fabian.arellano - 15:37, Friday 22 May 2020 (14415)

Today I continued with the measurements.

Pending review of the quality of the data which may suggest repeating a measurement, I have finished the measurements with the control system on.

I still need to do the measurements with the control system off.

MIF (General)
Print this report.
fabian.arellano - 18:10, Thursday 21 May 2020 (14412)Get code to link to this report
On site shift report

I moved PR3in order to recover the transmitted green light in the X arm.

P: -6.8urad --> -9.7 urad

Y: 59 urad --> 60.7 urad

Transmitted green light:

Y arm: 0.72 (I didn"t do anything)

X arm: 0.2 -->  0.8

IMC output power: 2.2 W

MIF (General)
Print this report.
masayuki.nakano - 16:13, Wednesday 20 May 2020 (14411)Get code to link to this report
Comment to PRFPMI Lock trial (Click here to view original report: 14409)

I aligned whole IFO again, and all value is within reasonable value (>0.9).

So far, nothing seems to have any problem, although the PRFPMI cannot be locked due to the violent ocean condition.

MIF (General)
Print this report.
shinji.miyoki - 10:25, Wednesday 20 May 2020 (14410)Get code to link to this report
Comment to PRFPMI Lock trial (Click here to view original report: 14409)


I guess OMMTs alignment were so shaked yesterday.

In addition to this, SR2 had troubles. So SR2 alignment should be healthy.

MIF (General)
Print this report.
takaaki.yokozawa - 08:36, Wednesday 20 May 2020 (14409)Get code to link to this report
PRFPMI Lock trial

[Yokozawa]

I tried to lock the PRFPMI today, but failed. I wrote the memo for lock trial.

1. SAL alignment worked (partially) well.

Xarm/Yarm IR lock with GR were succeeded, PRMI lock succeeded. But IR trans light(X~0.8, Y~0.8) and POP90I RF PD(~0.7) signal were slighly lower than reference.

2. Micro-seismic motion was large

3. We couldn't see the light in the AS gige camera and OMC QPD.

Comments related to this report
shinji.miyoki - 10:25, Wednesday 20 May 2020 (14410)


I guess OMMTs alignment were so shaked yesterday.

In addition to this, SR2 had troubles. So SR2 alignment should be healthy.

masayuki.nakano - 16:13, Wednesday 20 May 2020 (14411)

I aligned whole IFO again, and all value is within reasonable value (>0.9).

So far, nothing seems to have any problem, although the PRFPMI cannot be locked due to the violent ocean condition.

VIS (SR2)
Print this report.
tatsuki.washimi - 15:28, Tuesday 19 May 2020 (14408)Get code to link to this report
Comment to IP does not move towards the set point. (Click here to view original report: 14381)

I checked the other PEM channels around SR2.

No any strange large signals are detected.

Images attached to this comment
14408_20200519082841_20200519.png
VIS (SR2)
Print this report.
fabian.arellano - 14:38, Tuesday 19 May 2020 (14407)Get code to link to this report
Comment to IP does not move towards the set point. (Click here to view original report: 14381)

In case you don't see the picture attached please click here: entry 14401.

Further investigation about the origin of the event suggests the cause was not an earthquake. The description of the picture is as follows:

  • The SR2 event happened at 23:42 UTC on May 7th when there's no earthquake.
  • The seismometer in the BS area clearly shows a perturbation at 00:33 UTC on May 8th, but nothing at 23:42 UTC on the 7th.
  • The perturbation mesured by the seismometer affected the feedback in SR3 IP.
  • Such perturbtion was likely the earthquake which happened southeast of Honshu island. See this report.

This report is related to entry 14404, in which the BS actuation is also shown.

MIF (ITF Control)
Print this report.
masayuki.nakano - 14:37, Tuesday 19 May 2020 (14406)Get code to link to this report
Recovery from the earthquake
Almost all has been recovered.
  • IMC is locked
  • We performed remote Trump method, and succeeded. The ramptime for the kick was 0.5. PR3 has been aligned to GOODOPLEV. OSEM signal seems fine.
  • PR2 is off-centered but in the oplev range. We cannot connected to pico (smart plug has been disconnected…), so now YamaT is going to go in the mine and reset them.
  • Green resonance was observed.
VIS (General)
Print this report.
fabian.arellano - 14:35, Tuesday 19 May 2020 (14404)Get code to link to this report
About BS and SR2 IP actuation.

While investigating what is reported in entries 14381 and 14401 I realized the BS IP actuation showed a similar behavior to SR2. See the picture:

  • IP H1 actuation began increasing on the 5th of May at 21:35 UTC in a similar way as SR2 on the 7th of May UTC.
  • SRM and SR3 didn't show a similar behavior in the interval of time examined.
  • Just as a reminder, as pointed in one of the entries cited above, the SR2 event doesn't seem to be produced by an earthquake.

Fortunately the BS IP coil driver is still functional. More investigation is needed.

Images attached to this report
14404_20200519073427_bsandsr2events.png
PEM (Center)
Print this report.
takaaki.yokozawa - 14:34, Tuesday 19 May 2020 (14405)Get code to link to this report
Comment to Large earthquake at hida (Click here to view original report: 14402)

It still continues the earthquake in the Hida city.

I checked the vacuum and temperature from the request by Miyoki-san.

Now, I could not see any diffrence from normal operation(Except of the OMC vacuum monitor).

Images attached to this comment
14405_20200519073340_55.png 14405_20200519073354_23.png 14405_20200519073402_50.png 14405_20200519073410_09.png
PEM (Center)
Print this report.
osamu.miyakawa - 13:46, Tuesday 19 May 2020 (14403)Get code to link to this report
Comment to Large earthquake at hida (Click here to view original report: 14402)
2020/05/19:13h29m Nagano M2.5
2020/05/19:13h23m Hida    M3.0
2020/05/19:13h17m Hida    M3.2
2020/05/19:13h13m Hida    M5.3
2020/05/19:13h46m Hida    M3.7
2020/05/19:05h10m Nagano    M3.4
2020/05/19:05h00m Hida   M4.0
2020/05/19:04h26m Hida   M2.4
2020/05/19:03h50m Hida   M2.7
2020/05/19:03h04m Hida   M2.6
2020/05/19:02h57m Hida   M3.3
2020/05/19:02h36m Hida   M2.8
2020/05/19:02h32m Hida   M2.5
2020/05/19:02h30m Hida   M2.9
2020/05/19:02h03m Hida   M3.0
2020/05/19:02h01m Hida   M4.8
2020/05/15:18h09m Hida   M2.6
2020/05/15:01h57m Hida   M2.7
2020/05/15:00h24m Hida   M3.1
2020/05/14:10h58m Hida   M2.6
2020/05/14:09h54m Hida   M2.5
2020/05/14:09h39m Hida   M2.9
2020/05/14:00h29m Hida   M3.4
2020/05/13:22h20m Hida   M3.1
2020/05/13:15h15m Hida   M3.7
2020/05/13:11h06m Hida   M2.7
2020/05/13:10h28m Hida   M4.6
2020/05/13:09h40m Hida   M4.5
2020/05/13:09h16m Hida   M2.6
2020/05/13:08h19m Hida   M2.6
2020/05/13:07h31m Hida   M3.5
2020/05/13:07h21m Nagano    M2.8
2020/05/13:07h07m Hida   M3.2
2020/05/13:07h04m Hida   M4.5
2020/05/13:05h42m Hida   M2.9
2020/05/13:05h28m Nagano    M3.0
2020/05/13:03h44m Hida   M2.8
2020/05/19:13h29m Nagano M2.5
2020/05/19:13h23m Hida    M3.0
2020/05/19:13h17m Hida    M3.2
2020/05/19:13h13m Hida    M5.3
2020/05/19:13h46m Hida    M3.7
2020/05/19:05h10m Nagano    M3.4
2020/05/19:05h00m Hida   M4.0
2020/05/19:04h26m Hida   M2.4
2020/05/19:03h50m Hida   M2.7
2020/05/19:03h04m Hida   M2.6
2020/05/19:02h57m Hida   M3.3
2020/05/19:02h36m Hida   M2.8
2020/05/19:02h32m Hida   M2.5
2020/05/19:02h30m Hida   M2.9
2020/05/19:02h03m Hida   M3.0
2020/05/19:02h01m Hida   M4.8
2020/05/15:18h09m Hida   M2.6
2020/05/15:01h57m Hida   M2.7
2020/05/15:00h24m Hida   M3.1
2020/05/14:10h58m Hida   M2.6
2020/05/14:09h54m Hida   M2.5
2020/05/14:09h39m Hida   M2.9
2020/05/14:00h29m Hida   M3.4
2020/05/13:22h20m Hida   M3.1
2020/05/13:15h15m Hida   M3.7
2020/05/13:11h06m Hida   M2.7
2020/05/13:10h28m Hida   M4.6
2020/05/13:09h40m Hida   M4.5
2020/05/13:09h16m Hida   M2.6
2020/05/13:08h19m Hida   M2.6
2020/05/13:07h31m Hida   M3.5
2020/05/13:07h21m Nagano    M2.8
2020/05/13:07h07m Hida   M3.2
2020/05/13:07h04m Hida   M4.5
2020/05/13:05h42m Hida   M2.9
2020/05/13:05h28m Nagano    M3.0
2020/05/13:03h44m Hida   M2.8
 
2020/05/19:13h29m Nagano M2.5
2020/05/19:13h23m Hida    M3.0
2020/05/19:13h17m Hida    M3.2
2020/05/19:13h13m Hida    M5.3
2020/05/19:13h46m Hida    M3.7
2020/05/19:05h10m Nagano    M3.4
2020/05/19:05h00m Hida   M4.0
2020/05/19:04h26m Hida   M2.4
2020/05/19:03h50m Hida   M2.7
2020/05/19:03h04m Hida   M2.6
2020/05/19:02h57m Hida   M3.3
2020/05/19:02h36m Hida   M2.8
2020/05/19:02h32m Hida   M2.5
2020/05/19:02h30m Hida   M2.9
2020/05/19:02h03m Hida   M3.0
2020/05/19:02h01m Hida   M4.8
2020/05/15:18h09m Hida   M2.6
2020/05/15:01h57m Hida   M2.7
2020/05/15:00h24m Hida   M3.1
2020/05/14:10h58m Hida   M2.6
2020/05/14:09h54m Hida   M2.5
2020/05/14:09h39m Hida   M2.9
2020/05/14:00h29m Hida   M3.4
2020/05/13:22h20m Hida   M3.1
2020/05/13:15h15m Hida   M3.7
2020/05/13:11h06m Hida   M2.7
2020/05/13:10h28m Hida   M4.6
2020/05/13:09h40m Hida   M4.5
2020/05/13:09h16m Hida   M2.6
2020/05/13:08h19m Hida   M2.6
2020/05/13:07h31m Hida   M3.5
2020/05/13:07h21m Nagano    M2.8
2020/05/13:07h07m Hida   M3.2
2020/05/13:07h04m Hida   M4.5
2020/05/13:05h42m Hida   M2.9
2020/05/13:05h28m Nagano    M3.0
2020/05/13:03h44m Hida   M2.8
 
List of earthquakes in a  week.
There was no earthquake for a whille after 5/14, but we have many from this morning.
 
2020/05/19:13h29m Nagano M2.5
2020/05/19:13h23m Hida    M3.0
2020/05/19:13h17m Hida    M3.2
2020/05/19:13h13m Hida    M5.3
2020/05/19:13h46m Hida    M3.7
2020/05/19:05h10m Nagano    M3.4
2020/05/19:05h00m Hida   M4.0
2020/05/19:04h26m Hida   M2.4
2020/05/19:03h50m Hida   M2.7
2020/05/19:03h04m Hida   M2.6
2020/05/19:02h57m Hida   M3.3
2020/05/19:02h36m Hida   M2.8
2020/05/19:02h32m Hida   M2.5
2020/05/19:02h30m Hida   M2.9
2020/05/19:02h03m Hida   M3.0
2020/05/19:02h01m Hida   M4.8
2020/05/15:18h09m Hida   M2.6
2020/05/15:01h57m Hida   M2.7
2020/05/15:00h24m Hida   M3.1
2020/05/14:10h58m Hida   M2.6
2020/05/14:09h54m Hida   M2.5
2020/05/14:09h39m Hida   M2.9
2020/05/14:00h29m Hida   M3.4
2020/05/13:22h20m Hida   M3.1
2020/05/13:15h15m Hida   M3.7
2020/05/13:11h06m Hida   M2.7
2020/05/13:10h28m Hida   M4.6
2020/05/13:09h40m Hida   M4.5
2020/05/13:09h16m Hida   M2.6
2020/05/13:08h19m Hida   M2.6
2020/05/13:07h31m Hida   M3.5
2020/05/13:07h21m Nagano    M2.8
2020/05/13:07h07m Hida   M3.2
2020/05/13:07h04m Hida   M4.5
2020/05/13:05h42m Hida   M2.9
2020/05/13:05h28m Nagano    M3.0
2020/05/13:03h44m Hida   M2.8
 
VIS (SR2)
Print this report.
fabian.arellano - 13:45, Tuesday 19 May 2020 (14401)Get code to link to this report
Comment to IP does not move towards the set point. (Click here to view original report: 14381)

In case you don't see the picture attached please click here: entry 14401.

Further investigation about the origin of the event suggests the cause was not an earthquake. The description of the picture is as follows:

  • The SR2 event happened at 23:42 UTC on May 7th when there's no earthquake.
  • The seismometer in the BS area clearly shows a perturbation at 00:33 UTC on May 8th, but nothing at 23:42 UTC on the 7th.
  • The perturbation mesured by the seismometer affected the feedback in SR3 IP.
  • Such perturbtion was likely the earthquake which happened southeast of Honshu island. See this report.

There will be a report in the VIS General section about this event, please stay tuned.

Images attached to this comment
14401_20200519064425_sr2eventmay8thjst.png
PEM (Center)
Print this report.
takaaki.yokozawa - 13:29, Tuesday 19 May 2020 (14402)Get code to link to this report
Large earthquake at hida

Magnitude 5.3 earthquake happend around hida-area.

We need the health check for them.

Images attached to this report
14402_20200519062923_42.png
Comments related to this report
osamu.miyakawa - 13:46, Tuesday 19 May 2020 (14403)
2020/05/19:13h29m Nagano M2.5
2020/05/19:13h23m Hida    M3.0
2020/05/19:13h17m Hida    M3.2
2020/05/19:13h13m Hida    M5.3
2020/05/19:13h46m Hida    M3.7
2020/05/19:05h10m Nagano    M3.4
2020/05/19:05h00m Hida   M4.0
2020/05/19:04h26m Hida   M2.4
2020/05/19:03h50m Hida   M2.7
2020/05/19:03h04m Hida   M2.6
2020/05/19:02h57m Hida   M3.3
2020/05/19:02h36m Hida   M2.8
2020/05/19:02h32m Hida   M2.5
2020/05/19:02h30m Hida   M2.9
2020/05/19:02h03m Hida   M3.0
2020/05/19:02h01m Hida   M4.8
2020/05/15:18h09m Hida   M2.6
2020/05/15:01h57m Hida   M2.7
2020/05/15:00h24m Hida   M3.1
2020/05/14:10h58m Hida   M2.6
2020/05/14:09h54m Hida   M2.5
2020/05/14:09h39m Hida   M2.9
2020/05/14:00h29m Hida   M3.4
2020/05/13:22h20m Hida   M3.1
2020/05/13:15h15m Hida   M3.7
2020/05/13:11h06m Hida   M2.7
2020/05/13:10h28m Hida   M4.6
2020/05/13:09h40m Hida   M4.5
2020/05/13:09h16m Hida   M2.6
2020/05/13:08h19m Hida   M2.6
2020/05/13:07h31m Hida   M3.5
2020/05/13:07h21m Nagano    M2.8
2020/05/13:07h07m Hida   M3.2
2020/05/13:07h04m Hida   M4.5
2020/05/13:05h42m Hida   M2.9
2020/05/13:05h28m Nagano    M3.0
2020/05/13:03h44m Hida   M2.8
2020/05/19:13h29m Nagano M2.5
2020/05/19:13h23m Hida    M3.0
2020/05/19:13h17m Hida    M3.2
2020/05/19:13h13m Hida    M5.3
2020/05/19:13h46m Hida    M3.7
2020/05/19:05h10m Nagano    M3.4
2020/05/19:05h00m Hida   M4.0
2020/05/19:04h26m Hida   M2.4
2020/05/19:03h50m Hida   M2.7
2020/05/19:03h04m Hida   M2.6
2020/05/19:02h57m Hida   M3.3
2020/05/19:02h36m Hida   M2.8
2020/05/19:02h32m Hida   M2.5
2020/05/19:02h30m Hida   M2.9
2020/05/19:02h03m Hida   M3.0
2020/05/19:02h01m Hida   M4.8
2020/05/15:18h09m Hida   M2.6
2020/05/15:01h57m Hida   M2.7
2020/05/15:00h24m Hida   M3.1
2020/05/14:10h58m Hida   M2.6
2020/05/14:09h54m Hida   M2.5
2020/05/14:09h39m Hida   M2.9
2020/05/14:00h29m Hida   M3.4
2020/05/13:22h20m Hida   M3.1
2020/05/13:15h15m Hida   M3.7
2020/05/13:11h06m Hida   M2.7
2020/05/13:10h28m Hida   M4.6
2020/05/13:09h40m Hida   M4.5
2020/05/13:09h16m Hida   M2.6
2020/05/13:08h19m Hida   M2.6
2020/05/13:07h31m Hida   M3.5
2020/05/13:07h21m Nagano    M2.8
2020/05/13:07h07m Hida   M3.2
2020/05/13:07h04m Hida   M4.5
2020/05/13:05h42m Hida   M2.9
2020/05/13:05h28m Nagano    M3.0
2020/05/13:03h44m Hida   M2.8
 
2020/05/19:13h29m Nagano M2.5
2020/05/19:13h23m Hida    M3.0
2020/05/19:13h17m Hida    M3.2
2020/05/19:13h13m Hida    M5.3
2020/05/19:13h46m Hida    M3.7
2020/05/19:05h10m Nagano    M3.4
2020/05/19:05h00m Hida   M4.0
2020/05/19:04h26m Hida   M2.4
2020/05/19:03h50m Hida   M2.7
2020/05/19:03h04m Hida   M2.6
2020/05/19:02h57m Hida   M3.3
2020/05/19:02h36m Hida   M2.8
2020/05/19:02h32m Hida   M2.5
2020/05/19:02h30m Hida   M2.9
2020/05/19:02h03m Hida   M3.0
2020/05/19:02h01m Hida   M4.8
2020/05/15:18h09m Hida   M2.6
2020/05/15:01h57m Hida   M2.7
2020/05/15:00h24m Hida   M3.1
2020/05/14:10h58m Hida   M2.6
2020/05/14:09h54m Hida   M2.5
2020/05/14:09h39m Hida   M2.9
2020/05/14:00h29m Hida   M3.4
2020/05/13:22h20m Hida   M3.1
2020/05/13:15h15m Hida   M3.7
2020/05/13:11h06m Hida   M2.7
2020/05/13:10h28m Hida   M4.6
2020/05/13:09h40m Hida   M4.5
2020/05/13:09h16m Hida   M2.6
2020/05/13:08h19m Hida   M2.6
2020/05/13:07h31m Hida   M3.5
2020/05/13:07h21m Nagano    M2.8
2020/05/13:07h07m Hida   M3.2
2020/05/13:07h04m Hida   M4.5
2020/05/13:05h42m Hida   M2.9
2020/05/13:05h28m Nagano    M3.0
2020/05/13:03h44m Hida   M2.8
 
List of earthquakes in a  week.
There was no earthquake for a whille after 5/14, but we have many from this morning.
 
2020/05/19:13h29m Nagano M2.5
2020/05/19:13h23m Hida    M3.0
2020/05/19:13h17m Hida    M3.2
2020/05/19:13h13m Hida    M5.3
2020/05/19:13h46m Hida    M3.7
2020/05/19:05h10m Nagano    M3.4
2020/05/19:05h00m Hida   M4.0
2020/05/19:04h26m Hida   M2.4
2020/05/19:03h50m Hida   M2.7
2020/05/19:03h04m Hida   M2.6
2020/05/19:02h57m Hida   M3.3
2020/05/19:02h36m Hida   M2.8
2020/05/19:02h32m Hida   M2.5
2020/05/19:02h30m Hida   M2.9
2020/05/19:02h03m Hida   M3.0
2020/05/19:02h01m Hida   M4.8
2020/05/15:18h09m Hida   M2.6
2020/05/15:01h57m Hida   M2.7
2020/05/15:00h24m Hida   M3.1
2020/05/14:10h58m Hida   M2.6
2020/05/14:09h54m Hida   M2.5
2020/05/14:09h39m Hida   M2.9
2020/05/14:00h29m Hida   M3.4
2020/05/13:22h20m Hida   M3.1
2020/05/13:15h15m Hida   M3.7
2020/05/13:11h06m Hida   M2.7
2020/05/13:10h28m Hida   M4.6
2020/05/13:09h40m Hida   M4.5
2020/05/13:09h16m Hida   M2.6
2020/05/13:08h19m Hida   M2.6
2020/05/13:07h31m Hida   M3.5
2020/05/13:07h21m Nagano    M2.8
2020/05/13:07h07m Hida   M3.2
2020/05/13:07h04m Hida   M4.5
2020/05/13:05h42m Hida   M2.9
2020/05/13:05h28m Nagano    M3.0
2020/05/13:03h44m Hida   M2.8
 
takaaki.yokozawa - 14:34, Tuesday 19 May 2020 (14405)

It still continues the earthquake in the Hida city.

I checked the vacuum and temperature from the request by Miyoki-san.

Now, I could not see any diffrence from normal operation(Except of the OMC vacuum monitor).